convert 5378.27 and 21.78 to Floating Point Standard (FPS) modified (16 bits) by changing 23 bit...

convert

5378.27 and 21.78 to Floating Point Standard (FPS) modified (16 bits) by changing 23 bit fractional part to a 7 bit fractional part. then multiply
Answer #1
Answer: 0 10001111 1100100

Explanation:
-------------
Converting 5378.27 to binary
   Convert decimal part first, then the fractional part
   > First convert 5378 to binary
   Divide 5378 successively by 2 until the quotient is 0
      > 5378/2 = 2689, remainder is 0
      > 2689/2 = 1344, remainder is 1
      > 1344/2 = 672, remainder is 0
      > 672/2 = 336, remainder is 0
      > 336/2 = 168, remainder is 0
      > 168/2 = 84, remainder is 0
      > 84/2 = 42, remainder is 0
      > 42/2 = 21, remainder is 0
      > 21/2 = 10, remainder is 1
      > 10/2 = 5, remainder is 0
      > 5/2 = 2, remainder is 1
      > 2/2 = 1, remainder is 0
      > 1/2 = 0, remainder is 1
   Read remainders from the bottom to top as 1010100000010
   So, 5378 of decimal is 1010100000010 in binary
   > Now, Convert 0.27000000 to binary
      > Multiply 0.27000000 with 2.  Since 0.54000000 is < 1. then add 0 to result
      > Multiply 0.54000000 with 2.  Since 1.08000000 is >= 1. then add 1 to result
      > Multiply 0.08000000 with 2.  Since 0.16000000 is < 1. then add 0 to result
      > Multiply 0.16000000 with 2.  Since 0.32000000 is < 1. then add 0 to result
      > Multiply 0.32000000 with 2.  Since 0.64000000 is < 1. then add 0 to result
      > Multiply 0.64000000 with 2.  Since 1.28000000 is >= 1. then add 1 to result
      > This is equal to 1, so, stop calculating
   0.27000000000043656 of decimal is .010001 in binary
   so, 5378.27 in binary is 1010100000010.010001
5378.27 in simple binary => 1010100000010.010001
so, 5378.27 in normal binary is 1010100000010.010001 => 1.0101 * 2^12

16-bit format:
--------------------
sign bit is 0(+ve)
exp bits are (127+12=139) => 10001011
   Divide 139 successively by 2 until the quotient is 0
      > 139/2 = 69, remainder is 1
      > 69/2 = 34, remainder is 1
      > 34/2 = 17, remainder is 0
      > 17/2 = 8, remainder is 1
      > 8/2 = 4, remainder is 0
      > 4/2 = 2, remainder is 0
      > 2/2 = 1, remainder is 0
      > 1/2 = 0, remainder is 1
   Read remainders from the bottom to top as 10001011
   So, 139 of decimal is 10001011 in binary
frac bits are 0101000

so, 5378.27 in 16-bit format is 0 10001011 0101000

Converting 21.78 to binary
   Convert decimal part first, then the fractional part
   > First convert 21 to binary
   Divide 21 successively by 2 until the quotient is 0
      > 21/2 = 10, remainder is 1
      > 10/2 = 5, remainder is 0
      > 5/2 = 2, remainder is 1
      > 2/2 = 1, remainder is 0
      > 1/2 = 0, remainder is 1
   Read remainders from the bottom to top as 10101
   So, 21 of decimal is 10101 in binary
   > Now, Convert 0.78000000 to binary
      > Multiply 0.78000000 with 2.  Since 1.56000000 is >= 1. then add 1 to result
      > Multiply 0.56000000 with 2.  Since 1.12000000 is >= 1. then add 1 to result
      > Multiply 0.12000000 with 2.  Since 0.24000000 is < 1. then add 0 to result
      > Multiply 0.24000000 with 2.  Since 0.48000000 is < 1. then add 0 to result
      > Multiply 0.48000000 with 2.  Since 0.96000000 is < 1. then add 0 to result
      > Multiply 0.96000000 with 2.  Since 1.92000000 is >= 1. then add 1 to result
      > This is equal to 1, so, stop calculating
   0.7800000000000011 of decimal is .110001 in binary
   so, 21.78 in binary is 10101.110001
21.78 in simple binary => 10101.110001
so, 21.78 in normal binary is 10101.110001 => 1.010111 * 2^4

16-bit format:
--------------------
sign bit is 0(+ve)
exp bits are (127+4=131) => 10000011
   Divide 131 successively by 2 until the quotient is 0
      > 131/2 = 65, remainder is 1
      > 65/2 = 32, remainder is 1
      > 32/2 = 16, remainder is 0
      > 16/2 = 8, remainder is 0
      > 8/2 = 4, remainder is 0
      > 4/2 = 2, remainder is 0
      > 2/2 = 1, remainder is 0
      > 1/2 = 0, remainder is 1
   Read remainders from the bottom to top as 10000011
   So, 131 of decimal is 10000011 in binary
frac bits are 0101110

so, 21.78 in 16-bit format is 0 10000011 0101110

5378.27 x 21.78 = 1.0101 * 2^12 x 1.010111 * 2^4 = 1.1100100011 * 2^16

1.1100100011 * 2^16

16-bit format:
--------------------
sign bit is 0(+ve)
exp bits are (127+16=143) => 10001111
   Divide 143 successively by 2 until the quotient is 0
      > 143/2 = 71, remainder is 1
      > 71/2 = 35, remainder is 1
      > 35/2 = 17, remainder is 1
      > 17/2 = 8, remainder is 1
      > 8/2 = 4, remainder is 0
      > 4/2 = 2, remainder is 0
      > 2/2 = 1, remainder is 0
      > 1/2 = 0, remainder is 1
   Read remainders from the bottom to top as 10001111
   So, 143 of decimal is 10001111 in binary
frac bits are 1100100

so, 5378.27 x 21.78 in 16-bit format is 0 10001111 1100100

Answer: 0 10001111 1100100